OpenStudy (anonymous):

could someone maybe help me out and check this?: use the formula v1xv2=|v1||v2|cos theta to find the cosine of the angle between the vectors v= (-7, 8) and w= (5, 3). I got -11/63.78-------> is that correct or am i way off?

OpenStudy (anonymous):

\[v\cdot w=(-7)(5)+(8)(3)=-35+24=-11\] \[|v|=\sqrt{(-7)^2+8^2}=\sqrt{113}\] \[|w|=\sqrt{5^2+3^2}=\sqrt{34}\] \[-11=\sqrt{3842}\cos\theta~~\iff~~\cos\theta\approx-\frac{11}{61.984}~~\iff~~\theta=\cdots\]

OpenStudy (anonymous):

would it be .99? @SithsAndGiggles